In a given normal distribution, the population mean is 80 and the population standard deviation is 2.5. Find the corresponding z-score of the following values:

15. 84


1
Expert's answer
2022-05-08T11:57:17-0400

a)


"Z=\dfrac{15-80}{2.5}=-26"

b)


"Z=\dfrac{84-80}{2.5}=1.6"

c)


"Z=\dfrac{15.84-80}{2.5}=-25.664"